tutor quick help plz and solution. One printing press can finish a certain job in 1 1/2 hours which a smaller press had been working on for half an hour when it broke down and the job had to be transferred to the smaller press. How long did it take the smaller press to finish the job?
Answer by richwmiller(17219)


Very confusing. Poorly written problem.
which press broke down?
Which press was working ?
I decided it must be the larger press broke down and both had been working
.5/1.5+.5/y=1
1/3+.5/y=1
.5/y=2/3
3*.5=2y
1.5=2y
y=.75 =3/4 =45 minutes
check
.5/1.5=.5/.75=1
1/3+2/3=1
1=1
BTW A good way to be ignored is ask for quick help. It generally means you are waiting to the last minute to do your homework and are just going to copy and hand it in as your own..